OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). sec2x - 2 = tan2x

OpenStudy (raden):

(sec2x - 2 = tan2x) or (sec^2 x - 2 = tan^2 x) ?

OpenStudy (anonymous):

\[\sec^2x-2 +\tan^2x\]

OpenStudy (raden):

= 0 right ?

OpenStudy (anonymous):

\[\sec^2x−2=\tan^2x\] sorry

OpenStudy (raden):

looks your equation still wrong if i use the identity : sec^2 x = tan^2 x + 1 your equation can be sec^2 x - 2 = tan^2 x tan^2 x + 1 - 2 = tan^2 x -1 = 0 nah, recheck again ur equ, please

OpenStudy (anonymous):

sec^2 x - 2 = tan^2 x

OpenStudy (dumbcow):

yeah in the case where you have a contradiction ...1=0 the answer is no solution use @RadEn explanation